OpenStudy (anonymous):

7p+6d-28=50 whats the value of p and d ???!!!!

OpenStudy (ash2326):

We need another equation to find p and d!!! Is there any other equation?

OpenStudy (anonymous):

7p + 6d - 28 = 50 50 + 28 = 78 7 + 6 = 13 78 / 13 = 6 6 x 7 = 42 ( p=42) 6 x 6 = 36 ( d = 36) that's the way i would do it.. not 100% sure.. let me know

OpenStudy (anonymous):

good thanks

OpenStudy (anonymous):

yw